(a)Find the area of the region enclosed by the parabola 𝑦 = 2π‘₯ βˆ’ π‘₯

2 and the π‘₯ axis.

(b) Find the value of π‘š so that the line 𝑦 = π‘šπ‘₯ divides the region in part (a) into two

regions of equal area.


1
Expert's answer
2021-11-15T08:43:57-0500

(a) "\smallint_{0}^2 2x-x^2dx"


"=[x^2-\frac{x^3}{3}]_{0}^2"


"=4-\frac{8}{3}"


"=\frac{4}{3}"




(b)

calculate half of the area found in part A

"=\frac{4}{3}(\frac{1}{2})
\\=\frac{2}{3}"

For the function y = mx, one of the points is the origin (0,0). we need to find the other point that intersects the curve. proceed as follows

"mx=2x-x^2\\x^2+mx-2x=0\\x^2+x(m-2)=0
\\x=0 \ or \ x=2-m"


find x value in terms of m where curves intersect

"\smallint_{0}^{2-m}(2x-x^2-mx)dx=\frac{2}{3}"


"[x^2-\frac{x^3}{3}-\frac{mx^2}{2}]_{0}^{2-m}"


"(2-m)^2-\frac{(2-m)^3}{3}-\frac{m(2-m)^3}{2}=\frac{2}{3}"


"(2-m)^2[\frac{3}{3}-\frac{2}{3}+\frac{2m}{6}-\frac{3m}{6}=\frac{2}{3}"


"\frac{1}{6}(2-m)^2[2-m]=\frac{2}{3}"


"\frac{1}{6}(2-m)^3=\frac{2}{3}"

Evaluate and solve m

"(2-m)^3=\frac{2}{3}(6)"


"(2-m)=\sqrt[3]{4}"


"m=2-\sqrt[3]{4}"